Becton Labs, Inc., produces various chemical compounds for industrial use. One compound, called Fludex, is prepared using an elaborate distilling process. The company has developed standard costs for one unit of Fludex, as follows:

Standard Quantity
or Hours
Standard Price
or Rate
Standard Cost
Direct materials 2.10 ounces $ 24.00 per ounce $ 50.40
Direct labor 0.40 hours $ 14.00 per hour 5.60
Variable manufacturing overhead 0.40 hours $ 2.50 per hour 1.00
Total standard cost per unit $ 57.00

During November, the following activity was recorded related to the production of Fludex:

  1. Materials purchased, 11,500 ounces at a cost of $259,325.
  2. There was no beginning inventory of materials; however, at the end of the month, 2,700 ounces of material remained in ending inventory.

  3. The company employs 23 lab technicians to work on the production of Fludex. During November, they each worked an average of 100 hours at an average pay rate of $11.00 per hour.

  4. Variable manufacturing overhead is assigned to Fludex on the basis of direct labor-hours. Variable manufacturing overhead costs during November totaled $2,300.

  5. During November, the company produced 4,100 units of Fludex.

Required:

1. For direct materials:

a. Compute the price and quantity variances.

b. The materials were purchased from a new supplier who is anxious to enter into a long-term purchase contract. Would you recommend that the company sign the contract?

2. For direct labor:

a. Compute the rate and efficiency variances.

b. In the past, the 23 technicians employed in the production of Fludex consisted of 4 senior technicians and 19 assistants. During November, the company experimented with fewer senior technicians and more assistants in order to reduce labor costs. Would you recommend that the new labor mix be continued?

3. Compute the variable overhead rate and efficiency variances.
Answer #1

1.

Direct material price variance = Actual quantity*Standard price - Actual quantity*Actual price

Direct material price variance = 11,500*$24 - $259,325

Direct material price variance = $276,000 - 259,325 = $16,675 Favorable

Direct material quantity variance = Standard quantity*Standard price - Actual quantity*Standard price

Direct material quantity variance = 4,100*2.10*$24 - (11,500-2,700)*$24

Direct material quantity variance = $206,640 - 211,200 = $4,560 Unfavorable

b. I would recommend the company to sign the contract as Total direct material cost variance is favorable $12,115 ($16,675-4,560)

2.

Direct labor rate variance = Actual hours*Standard rate - Actual hours*Actual rate

Direct labor rate variance = 23*100*$14 - 23*100*$11

Direct labor rate variance = $32,200 - 25,300 = $6,900 Favorable

Direct labor efficiency variance = Standard hours*Standard rate - Actual hours*Actual rate

Direct labor efficiency variance = 4,100*0.40*$14 - 23*100*$14

Direct labor efficiency variance = $22,960 - 32,200 = $9,240 Unfavorable

b. I would recommend the company not to continue with the labor mix as total direct labor cost variance is unfavorable $2,340 ($6,900-9,240).

3.

Variable overhead rate variance = Actual hours*Standard rate - Actual hours*Actual rate

Variable overhead rate variance = 23*100*$2.50 - $2,300 = $3,450 Favorable

Variable overhead efficiency variance = Standard hours*Standard rate - Actual hours*Actual rate

Variable overhead efficiency variance = 4,100*0.40*$2.50 - 23*100*$2.50

Variable overhead efficiency variance = $4,100 - 5,750 = $1,650 Unfavorable